OpenStudy (jpes2193):

The force (f) of the wind on a flat surface positioned at a right angle to the direction of the wind varies jointly with the area A of the surface and the square of the speed of (v) of the wind. A wind of 40mph blowing on a window measuring 3 feet by 4 feet has a force of 160lbs. what is the force on the window measuring 4 feet by 5 feet caused by a wind of 90mph?

OpenStudy (jpes2193):

I understand the F=kxy the areas are 12 feet for the first windown and 20 feet for the second window. other than that im clueless on where to start with this problem.

OpenStudy (jpes2193):

@ribhu

OpenStudy (ribhu):

F= k (xy)(v)^2 160=k(12)(160) k=1/12

OpenStudy (ribhu):

now new case F = 1/120(20)(8100) F=8100/6 = 1350 lbs

OpenStudy (jpes2193):

what is k related to ?

OpenStudy (ribhu):

i think this should be the case. k is any thing which will homogenise this equation

OpenStudy (ribhu):

knowing k is not our concern here

OpenStudy (jpes2193):

okay so my teacher never taught us this... can you explain that differently ? and 1350 is one of the answers

OpenStudy (ribhu):